OpenStudy (anonymous):

Graph the function y=sqrt x-2-5 which point lies on the graph? a. (-5,2) b. (-2,5) c. (3,-4) d. (1,4)

OpenStudy (greatlife44):

@emilys819 whenever you have two functions say \[f*(g(x)) \] this means that you have two composite functions. If I remember this correctly, this means that you compute the value of x for whatever function is in the inside, then you compute the value for the other function. Let's look at this first: \[f(*g(x))= outside (inside) \] You plug in a value x in whatever function is on the inside and you'll get a number. you take that number and plug it into the function on the outside say f(x). first we identify our functions: f(x) = -4x+7 g(x) = 2x-6 f*(g(x)) in this case our x value is one, so we plug in one to g(x) and we will get a number. g(1) = 2(1)-6 = 2-4 = -4 so g(1) gives us -4 as our value. we take this and we plug this into f(x) now f(4) = -4(1)+7 = -4+7 = 3 so our answer for this is 3.

OpenStudy (greatlife44):

the answer is not 3, ah no, wait sorry.. we had -4 from the f(1) we take that number -4 and plug it into g(x) g(-4) -4(-4)+7 = 16+7 = 23

OpenStudy (anonymous):

thankyou!! def helped!!